OpenStudy (anonymous):

A 12 by 25 by 36 cm cereal box is lying on the floor on one of its 25 by 36 cm faces. An ant, located at one of the bottom corners of the box, must crawl along the outside of the box to reach the opposite bottom corner. What is the length of the shortest such path? Note: The ant can walk on any of the five faces of the box, except for the bottom face, which is flush in contact with the floor. It can crawl along any of the edges. It cannot crawl under the box. A 12 by 25 by 36 cm cereal box is lying on the floor on one of its 25 by 36 cm faces. An ant, located at one of the bottom corners of the box, must crawl along the outside of the box to reach the opposite bottom corner. What is the length of the shortest such path? Note: The ant can walk on any of the five faces of the box, except for the bottom face, which is flush in contact with the floor. It can crawl along any of the edges. It cannot crawl under the box. @Mathematics

OpenStudy (anonymous):

\[\sqrt{3697}=60.803 \text{ cm} \]Set the derivative of the following expression, the ant's total path length,\[2\left(\sqrt{x^2+12^2}+\sqrt{\left(\frac{25}{2}\right)^2+(18-x)^2}\right) \]to zero and solve for x.\[x=\frac{432}{49} \]Plug the value of x into the expression above and simplify